OpenStudy (calculusxy):

Help in the process of graphing y = x^2 + 5x + 2

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

what kind of curve is this ? any idea ?

OpenStudy (calculusxy):

Parabola

OpenStudy (phi):

yes. any idea what the x value of the vertex is ? a x^2 + bx + c x value of vertex is at -b/(2a)

OpenStudy (calculusxy):

\[\frac{ -b }{ 2a } = \frac{ -5 }{ 2(1) } = \frac{ -5 }{ 2} = -2.5 \]

OpenStudy (phi):

ok, and you can find the corresponding y value, right ?

OpenStudy (calculusxy):

yes \[y= -2.5^2 + 5(-2.5) + 2 = -4.25 \]

OpenStudy (phi):

ok, so the vertex is at (-2.5, -4.25) next, is this a smile or frown shape ?

OpenStudy (calculusxy):

smile?

OpenStudy (phi):

yes. You know because as you move away from x= 2.5 the parabola gets bigger (goes up) \( \cup\) shape

OpenStudy (phi):

we should find a few points. at x=0 y=2 (0,2) at x=1 : y= 1+5*1+2= 7 and a few points to the left of -2.5 say at x= -3 and x= -4

OpenStudy (phi):

we could find where y is 0, but we will get ugly numbers (we would use the quadratic formula to solve x^2 + 5x + 2 =0 (or complete the square) but it seems easier to find a few nice points, plot them, and sketch in the curve.

OpenStudy (phi):

hmmm, I should add correctly at x=1 : y= 1+5*1+2= 8

OpenStudy (calculusxy):

if i plot x = -3 then whatever get will be reflected wth -1 ?

OpenStudy (phi):

I am not sure what you are asking. but the y value at x=-3 will also be the y value "on the other side" of the vertex the vertex is at -2.5 -3 is 0.5 to the left, so you will also know the y value 0.5 to the right, at x= -2

OpenStudy (calculusxy):

i got it thanks!

OpenStudy (phi):

Here is a graph (from geogebra) with some of the points: you can see the symmetry that helps to plot other points.